OpenStudy (anonymous):

3 is subtracted from a number and then the difference is divided by 11. The result is 12. What is the number? (WORK BACKWARDS)

OpenStudy (zale101):

try working that out, and i'll correct u

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

n-3=n/11

OpenStudy (anonymous):

I honestly really don't get this question :/

OpenStudy (zale101):



OpenStudy (anonymous):

Wait I am confused. Is the 2 a 12? Or is this the formula or answer?

OpenStudy (zale101):

3 is subtracted from a number and then the difference is divided by 11. The result is 12. let's let the number be n for the variable 3 is subtracted from a number 3-n and then the difference is divided by 11 (3-n)/11 The result is 12. (3-n)/11=12

OpenStudy (zale101):